OpenStudy (anonymous):

find the equation of the circle passing through the point (-6,5), (-3,-4) and (2,1) what are the co-ordinates of the center of the circle?

OpenStudy (anonymous):

consider the gen eqn of the circle x^2+y^2+2gx+2fy+c=0, where g,f and c are constants to find out, put the points in this eq u will get three eqs in three unknows

OpenStudy (anonymous):

ok ... thanks ..

OpenStudy (anonymous):

can anyone give more explaination?

OpenStudy (anonymous):

did u form the three eqs?

OpenStudy (anonymous):

what r the values of g f and c ?

OpenStudy (anonymous):

for(-6,5) 36+25-12g+10f+c=0 -12g+10f+c=-61

OpenStudy (anonymous):

these r to determine by solving the eqs

OpenStudy (anonymous):

for (-3,-4) 9+16-6g-8f+c=0 -6g-8f+c=-25

OpenStudy (anonymous):

for (2,1) 4+1+4g+2f+c=0 4g+2f+c=-5

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

solve these eqs simultaneously to get g,f ,c

OpenStudy (anonymous):

then centre is (-g,-f) radius is sqrt(g^2+f^2-c)